OpenStudy (anonymous):

hello there, i'm Jastine Cabuyao. Mathematics student will you please help me.. we all know that polar coordinates is a two-dimensional coordinate system in which each point on a plane is determined by a distance from a fixed point and an angle from a fixed direction. will you please explain it to me further?

OpenStudy (mertsj):

A polar coordinate system uses the positive x axis as the ":pole" and a rotation in a counterclockwise direction from the pole to arrive at the fixed point. For example: (3, pi/4) means go 3 units to the right and rotate counterclockwise 45 degrees or pi/4.

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

there is a spider near you that is 36 degrees from your left, and 7 meters out ....

OpenStudy (anonymous):

polar co ordinates are the angular and radial representation of a vector. E.g a vector will have the polar co ordinates as r1=xcos(theta) and r2=ysin(theta)

OpenStudy (anonymous):

in finding the area of polar coordinates we use this formula \[A = \frac{ 1 }{ 2 } \int\limits_{\alpha}^{\beta} ( [f(\theta)]^2 - [g(\theta)] ^2 ) d \theta \]

OpenStudy (anonymous):

galing! :D

OpenStudy (anonymous):

echosera ka shedy hahaha alin ipapasa mo kay mam c?

OpenStudy (anonymous):

hahaha :D alin?na ganito ba? meron na ko :)